OpenStudy (anonymous):

Solve for x: 2x+1/ 3 = x+6/ 2

OpenStudy (anonymous):

familiar of cross multiplying?

OpenStudy (anonymous):

um i dont think so x.x

OpenStudy (anonymous):

if i have A C - = - B D what do yo think cross multiplying refer to?

OpenStudy (anonymous):

you would multiply a*d & b*c?

OpenStudy (anonymous):

very good, and what do you do with them?

OpenStudy (anonymous):

you would multiply

OpenStudy (anonymous):

well you already said that. How do you think a*d is related to b*c?

OpenStudy (anonymous):

they have variables? i am cross multiplying and i got 2(2x+1)=4x+2 and 3(x+6)= 3x+18

OpenStudy (anonymous):

so now i put them together and solve for x normally right?

OpenStudy (anonymous):

verry good

OpenStudy (anonymous):

so the answer would be x=16 correct?

OpenStudy (anonymous):

ikd, how did you get that?

OpenStudy (anonymous):

4x+2= 3x+18 (subtract 3x from both sides) x+2=18 (subtract 2 from both sides) x=16

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

Thank you! (:

OpenStudy (anonymous):

yw